3/4x+5/6=5x-125/3
Please solve that in an order

Answer by rfer(16322)

3/4x-5x=-125/3-5/6
-4 1/4x=- 250/6-5/6
-17/4x=-255/6
x=(-255/6)/(-17/4)
x=-255/6*-4/17
x=1020/102
x=10